OpenStudy (anonymous):

A roller coaster starts at point y1=55m and only gets it to 37m at the second hill. The total distance is 520m and the wagon weights 1300 kg. How high is the average frictional force?

OpenStudy (anonymous):

There has been a loss of energy that has turned into frictional work: If F is frictional force: \[F \times d=F \times 520 =m \times g \times (55-37)=1300 \times 9.8 \times 18\] Then: \[F=1300 \times 9.8 \times 18 / 520\]

OpenStudy (anonymous):

Thank you !